If the coordinates of A and B be (1, 1) and (5, 7), then the equation of the perpendicular bisector of the line segment AB is

Grade:11

1 Answers

Arun
5 years ago
 
 
Dear student
 
slope of AB = 6/4 = 3/2
hence slope of a line perpendicular to AB = – 2/3
 
now mid point of AB = (3,4)
 
hence equation of perpendicular bisector =
 
y – 4 = – 2/3 (x – 3)
 
y = – 2/3 x + 2 + 4
 
3y + 2x = 18